### v1

# Understanding the Issue Types in Open Source Blockchain-based Software Projects with the Transformer-based BERTopic

Abstract
Blockchain-based software systems are increasingly deployed across diverse domains, yet a systematic understanding of their development challenges remains limited. This paper presents a large-scale empirical study of 497,742 issues mined from 1,209 open-source blockchain projects hosted on GitHub. Employing BERTopic, a transformer-based topic modeling technique, we identify 49 distinct issue topics and organize them hierarchically into 11 major subcategories. Our analysis reveals that both general software development issues and blockchain-specific concerns are nearly equally represented, with Wallet Management and UI Enhancement emerging as the most prominent topics. We further examine the temporal evolution of issue categories and resolution times, finding that Wallet issues not only dominate in frequency but also exhibit the longest resolution time. Conversely, Mechanisms issues are resolved significantly faster. Issue frequency surged after 2016 with the rise of Ethereum and decentralized applications, but declined after 2022. These findings enhance our understanding of blockchain software maintenance, informing the development of specialized tools and practices to improve robustness and maintainability.
Index Terms:
Blockchain, Github Issues, Topic Modeling, BERTopic, Resolution Time
I Introduction
Blockchain technology has emerged as a transformative innovation with the potential to revolutionize a wide range of sectors [1] by offering enhanced transparency, security, and immutability. Beyond its foundational role in powering cryptocurrencies, blockchain is increasingly being adopted across diverse domains such as supply chain management [2, 3], healthcare [4], and finance [5]. The rapid growth of the cryptocurrency market, which has surpassed a trillion-dollar valuation [6], has further accelerated interest and investment in blockchain-based solutions. This widespread adoption has led to a surge in the development of blockchain application projects, creating a vibrant and rapidly evolving ecosystem of blockchain software development. Developing high-quality and maintainable applications requires a clear understanding of the nature and characteristics of software issues [7, 8] encountered during development, testing, and maintenance. These characteristics include the distribution of issues across different components [9], their impact on the system, resolution times, and the correlation between various issue types. Gaining insights into these aspects can guide the prioritization of development and testing [10] efforts. For instance, if a significant number of issues arise from cryptographic security flaws, it would indicate the need for more specialized tools, deeper domain expertise, and a stronger emphasis on security-focused development practices. Such understanding enables more informed decision-making and contributes to the creation of more robust, secure, and maintainable applications.
Recent research has focused on identifying issues and bugs in blockchain-based systems, aiming to uncover the underlying characteristics and trends that developers encounter. However, these studies have either focused narrowly on bugs within a limited number of blockchain projects [11] or examined other types of development activities, such as commits and pull requests [12]. To the best of our knowledge, no prior research has undertaken a systematic, large-scale analysis of issue reports from blockchain-based applications on GitHub to comprehensively characterize the development challenges prevalent in this domain.
To address this gap, we collected 497,742 issues from a curated set of 1,209 open-source blockchain-based projects on GitHub and applied topic modeling techniques to uncover the key development challenges faced by practitioners in this domain. Through this large-scale empirical analysis, we aim to answer the following three research questions:
RQ1: What are the key issue categories in blockchain projects as identified through topic analysis?
By leveraging advanced topic modeling techniques using BERTopic on GitHub issues, followed by systematic manual annotation, we identified and categorized 49 distinct topics across the analyzed projects. Among these, the most dominant were Wallet Management and Connectivity (21.3%), UI Enhancement (10.6%), and Logging and Error Handling (9.2%).
RQ2: How do resolution times vary across different subcategories of issues in blockchain projects?
The resolution time analysis shows that Wallet issues take notably longer to resolve, whereas Mechanisms issues, such as those involving consensus protocols, the Network Nervous System (NNS), and neuron staking, are resolved more quickly.
RQ3: How have different types of issues evolved over time in terms of frequency within blockchain projects?
We examined the temporal trends in issue creation across key subcategories and identified a significant surge in reported issues beginning around 2016. This surge was particularly pronounced in categories such as Development & Maintenance, User Experience & Interaction, and Wallet.
To facilitate replication and extension, we publicly share our data and code  [^1] https://github.com/SQMLab/StudyOnBlockchainApps.
II Background & Related Works
Numerous studies have explored issues, particularly from GitHub, to understand developer concerns, issue trends, and resolution patterns. Tamanna et al. [13] conducted a large-scale empirical study of 118K GitHub issues from 34 runtime system repositories to categorize issue types and analyze issue management practices. Beyer and Pinzger [14] conducted a manual classification of Android development issues on Stack Overflow, uncovering recurring challenges faced by mobile developers. In addition to manual categorization, various automated approaches have been used to analyze issue-fixing times, particularly for bugs. For example, Lamkanfi et al. [15] observed skewed bug-fix time distributions in Eclipse and Mozilla, while Panjer [16] applied logistic regression to predict bug resolution times using Eclipse Bugzilla data. Giger et al. [17] advanced this line of work using decision-tree-based models, achieving improved precision and recall. More recently, Ardimento et al. [18] applied BERT-based models to predict bug-fixing durations, reporting accuracy up to 91%.
Topic modeling has been extensively applied to mine latent themes in developer communications, including GitHub, Stack Overflow, and issue trackers. Thung et al. [8] and Uddin et al. [9] used Latent Dirichlet Allocation (LDA) to identify 40 prevalent topics in bug reports and developer Q&A, respectively. Bangash et al. [19] explored how developer discussions evolve temporally in the machine learning domain on Stack Overflow, revealing trends and tagging challenges by identifying 50 topics. More recently, BERTopic has emerged as a state-of-the-art topic modeling method that combines transformer-based embeddings with clustering techniques to extract coherent and semantically meaningful topics. Grootendorst [20] demonstrated that BERTopic outperforms classical models in coherence and human evaluation. Egger et al. [21] found BERTopic particularly effective for modeling short and noisy texts, such as Twitter data, demonstrating its versatility.
Several recent studies have applied topic modeling with GitHub issue analysis to study blockchain-related software development. Vaccargiu et al. [22] utilized BERTopic on GitHub issues and comments to uncover sustainability-related themes in Ethereum developer discussions, focusing on the Go-Ethereum repository. In a related study, Vaccargiu and Tonelli [23] applied BERTopic to blockchain projects in the environmental sector, identifying core concerns such as carbon credit tracking and resource optimization. Wan et al. [11] manually classified 946 unique bugs across eight blockchain systems into ten categories using card sorting. They also investigated bug-fixing dynamics in blockchain systems, revealing that security bugs took the longest, while performance bugs often remained unresolved for extended periods.
Despite extensive work on issue analysis, topic modeling, and blockchain software, there remains a gap in applying advanced topic modeling techniques to systematically uncover and categorize development challenges in blockchain-related GitHub issues. Existing work typically targets general software domains or specific concerns, lacking a comprehensive, data-driven view of developer-reported issues in blockchain projects. This study fills that gap.
III Methodology
The methodology is organized into two principal phases: (i) data collection and preprocessing of GitHub repositories and issues, and (ii) topic modeling using BERTopic to uncover recurring topics. Each phase is described in detail below.
III-A Data Collection
GitHub is a key platform for software artifacts, hosting millions of repositories with rich metadata (e.g., pull requests, issues, commits), offering valuable empirical insight into large-scale software development practices [24]. We therefore selected GitHub as our primary source for blockchain-related repositories. Figure 1 outlines the data collection pipeline.
Figure 1: Steps in data collection process.
III-A1 Project Selection
Identifying blockchain-related open-source repositories on GitHub is challenging due to inconsistent labeling and keyword usage, leading to search inaccuracies [12]. To address this, we adopted a semi-automated, two-phase selection strategy.
First, for repositories created before January 2021, we used the curated dataset from Das et al.[12], based on a validated GHTorrent subset [25]. This dataset originally included 1,194 repositories filtered by cryptocurrency relevance, platform affiliation, and popularity. However, as their dataset represents a snapshot up to January 2021, many of these repositories are found to be inactive or archived. Applying additional activity-based filtering, we retained 218 active repositories for analysis.
Second, since GHTorrent was discontinued after 2021, to collect the new repositories, we conducted a systematic query using the PyGithub [^2] https://github.com/PyGithub/PyGithub Python library that uses GitHub REST API [^3] https://docs.github.com/en/rest. The keyword list used in the search was adapted from Das et al.’s methodology [12]:
  • (i)
    names of major platforms and cryptocurrencies (e.g., Bitcoin, Ethereum), extracted from CoinMarketCap [^4] https://coinmarketcap.com;
  • (ii)
    names of prominent permissioned blockchain frameworks (e.g., Hyperledger Fabric, Corda), which lack cryptocurrency yet represent significant enterprise use cases; and
  • (iii)
    blockchain-related terms (e.g., smart contracts, dApps, Solidity) to capture relevant tools and applications.
To further refine the keyword set, popular decentralized application (dApp) names were incorporated based on rankings and traffic data from The Dapp List [^5] https://thedapplist.com, a curated directory highlighting trending dApps across multiple blockchains, and DappRadar [^6] https://dappradar.com, a widely-used analytics platform that tracks dApp usage statistics across blockchain networks. Including these sources helped ensure coverage of emerging and actively maintained projects that might otherwise be overlooked through generic keyword searches alone.
Additional automated filtering was applied to reduce noise: repositories were excluded if they were empty, archived, inactive in the past year, or had fewer than five stars or forks. These criteria, based on empirical guidelines [12, 24], aimed to balance dataset inclusiveness with quality. Inactivity was used as a proxy for abandonment, while stars and forks indicated public interest and collaboration. After filtering, 6,833 repositories remained.
III-A2 Project Filtering
Mining GitHub repositories presents several well-documented challenges. As highlighted by Kalliamvakou et al. [24], many do not represent actual software projects but instead host tutorials, assignments, or code snippets. Consequently, we undertook a comprehensive manual verification process to refine the initially collected 6,833 repositories. Each repository was evaluated individually, and those meeting any of the following conditions were excluded: (i) not related to blockchain software development; (ii) a demo application; (iii) used primarily for educational purposes (e.g., tutorials, course assignments, or coding exercises); (iv) serving as a static data or resource storage.
In summary, we excluded repositories not representing genuine blockchain software systems, resulting in 991 high-confidence repositories after 80+ hours of manual validation by the first two authors. Combined with 218 repositories from Das et al. [12], this yielded a final dataset of 1,209 repositories.
III-A3 Issue Collection
Following the identification of the target repositories, we collected all associated issues using the GitHub REST API via the PyGithub library. Notably, we discovered that 185 repositories contained no reported issues. Overall, we collected a total of 511,560 issues, encompassing both open and closed issues, without filtering by type.
Upon conducting a manual inspection of the issue corpus, we observed that a substantial number of issues consisted of a single word, such as Develop, Dev, Development, Master, Test, Staging, Backup, Release, etc., or a version number. These issues do not provide any information on the topic of the issue, but rather create ambiguity and lack analytical value. Consequently, we excluded all single-word issues from the dataset. After this step, we retained 497,742 issues, which served as the basis for the subsequent stages of our study.
III-B Topic Modeling with BERTopic
To identify and analyze topics within issues of blockchain software repositories, we employed BERTopic, a modern topic modeling framework that combines transformer-based embeddings, dimensionality reduction, and hierarchical clustering [20]. Given its superior performance over traditional methods [21, 26, 27, 28, 29], BERTopic was selected for this task.
Figure 2: BERTopic [20] can be viewed as a sequence of 5 steps to create its topic representations. The first three steps are for creating the topics, and the last three steps are for representing and interpreting the topics.
BERTopic offers extensive customization, requiring iterative tuning. Starting with default settings, we observed over a thousand imbalanced topics—some overly broad, others too narrow. This motivated a trial-and-error approach to refine parameters and optimize topic coherence.
BERTopic begins by converting documents into embeddings using transformer models. Unlike traditional topic modeling, which requires extensive data preprocessing [30, 19], it requires minimal preprocessing. As noted in the official BERTopic documentation [^7] https://maartengr.github.io/BERTopic/faq.html?#should-i-preprocess-the-data, traditional preprocessing is unnecessary when using document embeddings, as every part of a document contributes to understanding its overall topic. But, it is recommended to clean the text of excessive noise, such as HTML tags, as these do not contain contextual information. Following these, we only removed markup tags, emojis, and URLs from the data.
For numerical representation, we selected the all-mpnet-base-v2 model from the SentenceTransformers library. This model balances state-of-the-art embedding quality with computational efficiency and has been shown to perform robustly across a variety of Natural Language Processing tasks [31].
We initially used concatenated issue titles and bodies for topic modeling, but this led to noisy, semantically diffuse clusters with thousands of topics. Analysis revealed that many bodies were auto-generated by GitHub Actions or CI/CD tools or template-based, offering little useful content—an issue also noted in prior studies [32, 33]. To improve topic quality and interpretability, we limited modeling to issue titles, which are typically concise and semantically rich.
After generating an embedding, the next step is to cluster the embeddings. To facilitate faster clustering, we applied Uniform Manifold Approximation and Projection (UMAP) for dimensionality reduction. This choice is justified by UMAP’s ability to preserve global semantic relationships while enhancing local cluster separability [34]. We retained 100 dimensions and set the neighborhood size to 25, balancing computational efficiency with contextual richness in high-volume data.
Clustering was performed using Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN), a density-based algorithm well-suited for discovering arbitrarily shaped clusters and ignoring outliers [35, 36]. A minimum cluster size of 1,000 was set to capture dominant issue themes and reduce noise, improving topic label clarity. Initial experiments showed that smaller values produced thousands of topics, hindering manual analysis. Although this threshold resulted in 268,755 outliers, we ignored the outliers, prioritizing prominent and recurring topics, excluding infrequent ones.
After topic generation, BERTopic refines topic representation through three steps: c-TF-IDF with CountVectorizer to reduce frequent terms, and a hybrid approach combining KeyBERT-inspired extraction with Maximal Marginal Relevance (MMR) to enhance relevance and diversity. This process helps in providing meaningful names to each topic.
Finally, the selected BERTopic model configuration, with automatic topic count selection and probability estimation, yielded 49 distinct topics—a similar number to the 50 and 40 topics identified by Bangash et al. [19] and Uddin et al. [9], respectively, as noted in Section II.
IV Approach, Analysis & Results
In this section, we answer three research questions (RQ) based on the 49 topics generated by BERTopic.
RQ1: What are the key issue categories in blockchain projects as identified through topic analysis?
Figure 3: 49 Topics are generated using BERTopic.
After generating the topics, the next critical step involved assigning semantically meaningful and representative labels to each topic to facilitate interpretation. This labeling was performed collaboratively by the first and second authors, both researchers with expertise in software engineering and testing, through an iterative process of discussion and refinement. For each topic, 30 highly representative issues were examined—ranked based on their probability values—to determine the most suitable label. The process continued until both researchers reached a consensus. Figure 3 illustrates the distribution of issue percentages across all 49 identified topics. The top five most prominent topics identified across all issues were: Wallet Management and Connectivity (21.3%), UI Enhancement (10.6%), Logging and Error handling (9.2%), CICD (4.6%), and RPC and Networking (4.5%). The high volume of UI-related issues likely reflects the complexity of blockchain technology and unintuitive interfaces. Even popular wallets pose challenges: new users struggle with complex interactions [37], while experienced users face unfamiliar metaphors and transaction flows [38]. Our analysis supports this, with frequent terms like icon, modal, and mobile in issues.
After completing the labeling, the topics were further grouped into higher-level subcategories and categories by merging semantically and conceptually related topics into a hierarchical structure, guided by domain knowledge and following practices observed in prior research [39, 40]. For example, topics such as Write, Fix, Update Tests and E2E Test were placed under the broader subcategory of Testing. This categorization resulted in two overarching topics: Software Development Issues (53.87%) comprising 4 primary subcategories, and Blockchain-Specific Issues (46.13%) consisting of 7 primary subcategories. The hierarchical structure of these topics is illustrated in Figure 4.
Figure 4: Hierarchy of topics with 11 major subcategories.
From the constructed hierarchy, we selected 11 major subcategories to guide our analysis for the subsequent research questions. We provide a high-level summary of the 11 major subcategories, each accompanied by a brief description.
Development & Maintenance. This is the largest subcategory under Software Development Issues category, comprising 18 topics and accounting for 20.97% of all issues. These issues reflect the day-to-day challenges developers face in keeping software systems reliable and maintainable—such as resolving build failures, managing conflicting dependencies, addressing flaky tests, refining code quality, and fixing functional bugs.
User Experience & Interaction. This is the second largest subcategory under Software Development Issues category, comprising 19.6% of all issues. It includes seven topics that address design inconsistencies, language localization, translation issues, and notification features aimed at improving overall user interaction.
Infrastructure. One major topic of this subcategory, Observability (24,278 issues; 10.6%), includes concerns related to Logging and Error Handling, Event Analytics, and tools like Sentry, focusing on actionable logs, error tracking, diagnostics, and analytics integration. Meanwhile, the Data Access Interfaces subcategory addresses query efficiency, endpoint design, and schema evolution—highlighting the foundation for performance monitoring, debugging, and system integration.
Documentation. Comprising a single topic, this subcategory accounts for 3,267 (1.43%) issues of the total. These issues often address missing or outdated instructions, unclear usage guides, API documentation gaps, and requests for improved onboarding materials.
Wallet. Encompassing six topics and representing 25.53% of all issues, this subcategory is divided into two low-level subcategories. Wallet Management (24.28%) covers discussions related to managing wallet ledgers, MetaMask Browser Extensions, Tor, and Privacy Preserving Wallets. Meanwhile, Wallet Address Management (1.25%) focuses on issues involving QR code, address copying, and clipboard operations.
Digital Asset. This subcategory accounts for 6.24% of issues, primarily covering Token Management (4.81%) and Non-Fungible Tokens (NFTs) (1.43%). Token Management issues typically involve minting, transferring, balance synchronization, and integration with multi-chain or custom tokens. NFT-related discussions focus on metadata rendering, display inconsistencies, and listing functionality
Networking. Most issues in this subcategory relate to RPC failures, latency, and sync errors—often involving timeouts, broken connections, and unreliable data retrieval—affecting responsiveness and reliability in network-dependent tasks.
Transaction. The Transaction subcategory represents 3.53% of all issues (8,084 in total), encompassing concerns around transaction fees (2.52%) and token swap functionality (1.01%). Common issues include miscalculated or unexpectedly high gas fees, failed or delayed transactions, and poor user feedback during execution.
Community Interaction. This subcategory encompasses three distinct topics, accounting for 2.86% of all issues. These topics include Leaderboard and Scores (0.60%), which focuses on reputation tracking, Fundraising (0.60%), which deals with community-driven financial support, and Voting Proposals (1.66%), which highlights governance mechanisms.
Security & Data Integrity. This subcategory accounts for 2.38% of all issues (5,463 total), focusing on core cryptographic operations in blockchain systems. It includes Cryptographic Signing (1.8%), covering digital signatures and authentication workflows, and Inscription Management (0.58%), addressing secure on-chain data encoding and handling.
Mechanisms. It encompasses two topics: Consensus Mechanism (0.68%), which covers the rules and algorithms by which network nodes agree on the blockchain state, and Network Nervous System (NNS) and Neuron Staking (0.45%), which addresses governance and staking operations within the network’s neuron-based voting system.
These results indicate that wallet-related functionality, which is unique to blockchain ecosystems, represents the largest proportion of issues, and the second largest subcategory consists of software development and maintenance issues.
Wallet Management and Connectivity, UI Enhancement, and Logging and Error Handling are the most prevalent topics in blockchain projects. In hierarchical terms, Development & Maintenance and Wallet are the dominant subcategories, with blockchain-specific and general development issues appearing in nearly equal measure.
Figure 5: Percentage of issues resolved within different timeframes across subcategories, showing consistent resolution patterns.
RQ2: How do resolution times vary across different subcategories of issues in blockchain projects?
To examine how resolution times vary across subcategories, we computed the minimum, maximum, mean, and median durations for each subcategory (Table I). These statistics reveal notable differences in resolution patterns across issue types. For example, some issues are closed within minutes of creation. Instances like MystenLabs/sui/issues/1180888https://github.com/MystenLabs/sui/issues/1180 and Vagabonds-Labs/cofiblocks/issues/100 demonstrate this behavior. Prior studies have reported such cases where developers write a patch, commit it to the version control system (VCS), and subsequently create and immediately close the corresponding issue [11, 8, 15].
To determine whether the observed differences in resolution times across subcategories are statistically significant, we applied the non-parametric Kruskal-Wallis H-test, suitable for non-normally distributed data. The test yielded a highly significant result: $H=3640.67$, $p=0.0$, confirming that resolution times differ significantly across subcategories.
The Wallet subcategory exhibits the longest maximum resolution time of 7.9 years (2,896 days), followed by categories like Infrastructure and Transaction, which also have long extremes. On the other hand, Mechanisms shows the shortest maximum resolution time at 2.8 years (1,022 days), though most subcategories range between 4 and 7 years. These findings are consistent with previous work, which attributes long resolution times to poor issue descriptions, unclear reproducibility, or shifting project priorities [11, 8, 15].
The mean resolution time across subcategories generally ranges from approximately 20 to 46 days. Mechanisms issues have the shortest mean resolution time (20 days), indicating quicker turnarounds, while the Wallet category again stands out with the longest average (46 days), suggesting more complex or resource-intensive problem domains. Other categories, such as Security & Data Integrity and Transaction, also report higher means (37–45 days).
Regarding median resolution time, values range from just 0.6 to 2.9 days, further highlighting how many issues are resolved swiftly. The shortest median is seen in Mechanisms (0.62 days), and the longest in Security & Data Integrity (2.93 days). This skew, with a small median but relatively large mean, suggests the presence of outliers—most issues are resolved quickly, but a few take significantly longer.
TABLE I: Resolution Time (in days) of Closed Issues by Subcategory, Sorted by Mean Resolution Time.
Category Min Max Mean Median
Mechanisms 0.000035 1022.12 19.81 0.62
Development & Maintenance 0.000012 2201.83 20.50 0.70
Documentation 0.0 1595.29 20.69 0.60
User Experience & Interaction 0.000012 2537.03 27.50 0.87
Infrastructure 0.0 2218.66 28.82 0.98
Community Interaction 0.000046 1935.54 32.00 1.21
Networking 0.000035 2319.35 32.34 0.96
Digital Asset 0.0 2057.94 32.55 1.14
Transaction 0.000035 2474.69 36.65 1.80
Security & Data Integrity 0.000035 2446.68 45.47 2.93
Wallet 0.000023 2896.19 45.95 2.02
The histogram in Figure 5 reinforces these findings. It shows that a large majority of issues, particularly in categories like Mechanisms, Documentation, and Development & Maintenance, are resolved within a month. Over 85% of issues in most subcategories meet this threshold, with exceptions like Community Interaction, Wallet, Security & Data Integrity, and Transaction. Approximately 4–15% are resolved between one month and one year, while 1–4% take over a year.
These findings illustrate that while many issues are resolved quickly, subcategories related to blockchain, especially Wallets, require significantly longer attention and resources.
Resolution times vary remarkably across issue subcategories: Wallet issues take the longest, while Mechanisms issues are resolved the fastest. Nonetheless, most issues (around 80%) across all subcategories are resolved within a month.
RQ3: How have different types of issues evolved over time in terms of frequency within blockchain projects?
To understand how issue categories evolved over time, we analyzed quarterly issue counts for each of the 11 main subcategories from 2012 to 2024. Figure 6 visualizes these trends, capturing the trajectory of reported issues across the blockchain ecosystem.
Our analysis reveals a statistically significant variation in issue frequencies across subcategories, as confirmed by a Kruskal-Wallis H-test: $H=93.67$, $p=9.97\times 10^{-16}$. This non-parametric result indicates that the differences in distributions among the categories are not due to random chance, underscoring the heterogeneity in issue evolution.
Figure 6: Trends in reported issues by subcategory, quarterly from 2012 to 2024.
A significant rise in issue frequency is seen across nearly all subcategories beginning in 2016. A similar trend has also been observed in previous studies on dApps [41, 42]. This sharp increase coincides with the launch and growing adoption of platforms like Ethereum, which officially released its Homestead version in March 2016. Ethereum introduced a new paradigm of decentralized applications (dApps) and smart contracts, significantly expanding the scope of blockchain software development. This growth spurred the creation of a vast number of blockchain projects, which in turn generated an increasing number of reported issues, particularly in categories like Development & Maintenance, User Experience & Interaction, and Wallet.
From 2023 onward, we observed a noticeable decline in issue frequency across most subcategories. This trend may stem from ecosystem maturation, better documentation, and reusable tools that reduce the need for extensive issue tracking, though further research is needed. Despite the decline, subcategories like Wallet and UI continue to exhibit high activity, reflecting their central role in user interaction and asset management.
Among the individual categories, Development & Maintenance and Wallet show the strongest upward trends, suggesting these areas have received increasing focus. User Experience & Interaction and Infrastructure also exhibit strong growth, although they plateau slightly below the top-performing subcategories. In contrast, Documentation and Mechanisms lag behind, maintaining relatively lower positions throughout the timeline, indicating either slower progress or less emphasis. Community Interaction, Digital Asset, and Networking show steady progress, while Security & Data Integrity and Transaction fall in the mid-range with slightly fluctuating but overall upward trends. Overall, the trends show broad and varied characteristics of different subcategories, with some leading the development trajectory more than others.
The frequency of issues surged after 2016, coinciding with the rise of Ethereum and dApps, and then declined after 2022. Wallet, the most frequently reported subcategory, consistently remained the most reported throughout the timeline.
V Discussion
This study provides a large-scale empirical analysis of blockchain software projects using GitHub issue mining and BERTopic modeling. Analysis identifies the most common issue topics in blockchain projects and shows that Development & Maintenance and Wallet-related issues are most prominent. The dominance and prolonged resolution of wallet issues reflect their inherent complexity. This signals a pressing need for specialized tools, standard frameworks, and secure design practices tailored to wallet functionalities.
Despite longer resolution times for Wallet and Security issues, around 80% of issues are resolved within a month, indicating that most are relatively manageable. This suggests practitioners should allocate resources based on issue complexity, while researchers have an opportunity to develop blockchain-specific tools to address more demanding issues.
Trend analysis shows that the issue frequency rose after 2016 with Ethereum and dApps, then declined after 2022. Wallet issues remained prominent, with persistent UI and Infrastructure concerns. These trends show how platform evolution shapes development, urging practitioners to adapt and researchers to pursue longitudinal studies on evolving practices and tools.
Overall, this work reinforces the notion that blockchain software engineering extends beyond conventional development practices, requiring tailored methodologies, domain-specific tooling, and a deeper understanding of decentralized environments. For the research community, the findings offer a foundation for future studies aimed at mitigating the unique challenges of blockchain systems. For practitioners—such as software company leadership—the findings suggest focusing resources on high-issue areas like Wallet Management and Connectivity. By recognizing these domain-specific needs, both communities can contribute to more secure, maintainable, and user-friendly blockchain applications.
V-A Limitations
A limitation of this study is its exclusive focus on open-source blockchain projects, which may not fully capture the breadth of development practices and challenges across the broader blockchain ecosystem. Another limitation stems from relying solely on issue titles for topic modeling, due to the common presence of auto-generated or repetitive templates in issue bodies. However, many issues include rich, developer-written content that could provide deeper insights.
V-B Future Work
Future research can expand on this study by comparing blockchain-based projects with non-blockchain projects to highlight the unique challenges of blockchain technology. Exploring distinct blockchain domains, such as cryptocurrency, healthcare, finance, supply chain, and smart contracts, would provide a more comprehensive view of industry-specific issues. Investigating closed-source blockchain projects could further enrich our understanding by contrasting them with open-source environments. Finally, the quality of generated topics could be improved by selectively incorporating informative issue bodies while using automated filtering to exclude templated or auto-generated content.
VI Conclusion
This study investigates the challenges in blockchain-based software applications by mining and analyzing roughly half a million GitHub issues from more than a thousand open-source projects. Using topic modeling with BERTopic followed by manual annotation, 49 distinct topics were identified, which were further organized into a hierarchical structure comprising 11 major subcategories. General software issues slightly outnumber blockchain-specific ones, with frequent concerns including logging, wallet management, and UI enhancement. Issue resolution times varied, with wallet-related issues taking longer. While issue activity peaked around 2016 and declined after 2022, wallet issues remained consistently significant. The results show that blockchain development shares challenges with general software engineering but also faces unique issues that require dedicated solutions and practices.